OpenStudy (lgbasallote):

the speed of a nerve impulse in the human body is about 100 m/s. If you accidentally stub your toe in the dark, estimate the time it takes the nerve impulse to travel to your brain translation please? lol

OpenStudy (shubhamsrg):

i can translate this in hindi..lol :D

OpenStudy (lgbasallote):

then you can translate it in english too!! *_*

OpenStudy (shubhamsrg):

@_@

OpenStudy (lgbasallote):

blue eyes too!

OpenStudy (shubhamsrg):

lol,.1 blue other black!! some eye defect maybe! ;)

OpenStudy (lgbasallote):

okay..let's concentrate on this one now..i wouldnt want to fill this thread with 100+ replies again and my question goes unanswered

OpenStudy (shubhamsrg):

hehehe..sorry i dont have the ans..i was only killing/passing time! :D

OpenStudy (anonymous):

Distance from brain to toe in meters is..?

OpenStudy (anonymous):

well, it takes a finite time for your brain to register that pain. So, given that this signal travels along the nerves at 100m/s & given that the distance between you toe & your brain is about 6 feet, calculate the time it takes that info to reach the brain.

OpenStudy (anonymous):

\[vt=d\] \[v=100 m/s\] estimate \(d\)

OpenStudy (shubhamsrg):

ohh..get it now..hmmn

OpenStudy (lgbasallote):

i see...so we make the distance into 6 feet?

OpenStudy (anonymous):

Your height, yes

OpenStudy (lgbasallote):

that was easier than i thought o.O lol

OpenStudy (anonymous):

if we were being really precise d = length of human + length of foot ftw

OpenStudy (anonymous):

yeah, approximately. I think they just want you to get a feel for the time, qualitatively.

OpenStudy (anonymous):

java, how long is a string?

OpenStudy (anonymous):

thats easy, its limitless :)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

another reference :D http://www.youtube.com/watch?v=JLIwuWNOWBk

OpenStudy (lgbasallote):

suppose im 1.68 meters tall do i do it \[time = \frac{1.68}{100} = 0.0168s\]

OpenStudy (anonymous):

where are you going with this? is the the origin of your alias ?

OpenStudy (anonymous):

nope, java. just a relevant reference.

OpenStudy (anonymous):

click the url and watch that

OpenStudy (anonymous):

yes, lgba.

OpenStudy (anonymous):

v = d / t t = d /v

OpenStudy (lgbasallote):

great!!!

OpenStudy (anonymous):

il watch it later cheers

OpenStudy (anonymous):

now do you know what 17 milliseconds feels like?

OpenStudy (anonymous):

but we can still feel that lag? do u know what i mean?